Determine whether the ordered pair is a solution of the inequality.
(6,16) 4y-5x>5

Answer by mizjoice(22)

The simple answer is yes. but I suspect you need a more involved answer, so here goes: the first thing to remember is that in an ordered pair, the first number is the 'x', and the second number is the 'y'. In your problem, this would mean that x=6, and y=16.
After that it becomes a matter of substituting the numbers for those variables.
your problem will look like this 4(16)-5(6)>5. 4*16=64, 5*6=30. What you end up with is 64-30>5, 64-30=34, 34>5. 34 is greater than 5.